OpenStudy (anonymous):

Solve by the substitution method please 7x+y=-13 -5x+y=17 what is the order pair?

OpenStudy (anonymous):

I have tried to work it out myself and this is what I got. -5x+y(-7x-13)=17 So I going with x=-35/7 which gave me -5 so that is my x but I can get my y= yet. I tried and gt 22/20 thats not right?????

OpenStudy (anonymous):

Your process seemed right, but I got x to be -5/2

OpenStudy (anonymous):

Also realize there is no y once you substitute it into that one equation, I think you just had a typo though.

OpenStudy (anonymous):

but -5/2 is 2.5

OpenStudy (anonymous):

the 22/20 is what I got

OpenStudy (anonymous):

for y

OpenStudy (anonymous):

For y I got 9/2. Now you can check these numbers if you plug them into one of the equations and both sides are equal.

OpenStudy (anonymous):

So can you help me with it again

OpenStudy (anonymous):

Yes. So it looks like you tried to solve the first equation for y to get y=-7x-13 Then it looked like you plugged it into the 2nd equation and should have gotten -5x+(-7x-13)=17

OpenStudy (anonymous):

All you need to do next is solve for x, plug it back into one of the equations and solve for y.

OpenStudy (anonymous):

so when I do that then I get 13y+17/7y+5 and the whole problem is negative with a neg.sign front understand

OpenStudy (anonymous):

I don't quite see what you are doing...-5x-7x-13=17 then -12x=30 then x=-5/2 so plug -5/2 into either of the equations to get y. So for example, -5(-5/2)+y=17 then (25/2)+y=17 and y=9/2

OpenStudy (anonymous):

ok so what is the order pair

OpenStudy (anonymous):

can you please give me the order pair to this question

OpenStudy (anonymous):

(x,y) so (-5/2,9/2) which is the same as (-2.5,4.5)

OpenStudy (anonymous):

so that is the order pair?

OpenStudy (anonymous):

Yes. If you plug in both the x and y into either of those equations you should get the left side equal to the right.

OpenStudy (anonymous):

thanks again, is there a program that helps you work these out and learn them

OpenStudy (anonymous):

Not sure...

OpenStudy (anonymous):

so I need to put (-2.5,4.5) in or (-5/2, 9/2) in

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

when I put in the first one it said incorrect!!!

OpenStudy (anonymous):

The answers are correct based on the equations you gave me. I promise.

OpenStudy (anonymous):

sorry the top one is suppose to be 4y sorry so 7x+4y=-13 and -5x+y=17 I am so sorry I could not figure out what I was doing wrong

OpenStudy (anonymous):

Lol no problem, but that does change things :P

OpenStudy (anonymous):

http://www.webmath.com/solver2.html here is something I found that does solve systems of equations, well at least 2 of them. I looked over their work, and it seems like it can be confusing. I would recommend trying to do them on your own solving for x or y like we did, and maybe using that website to check steps if you made a mistake. There's probably something better out there somewhere.

OpenStudy (anonymous):

So can I get the order pair for this then please and I can move on to the next.....please...

OpenStudy (anonymous):

x=-3 y=2 so (-3,2)

OpenStudy (anonymous):

i have to continue working on ths because I am still a little confused but trying to understand. Thanks for the help again....I have more to come....

OpenStudy (anonymous):

No problem.